LASER INDUCED FLUORESCENCE SPECTROSCOPY SiCN : ROTATIONAL ANALYSIS OF THE Ã ^{2}Δ – X̃ ^{2}Π TRANSITION

\underline{MASARU FUKUSHIMA} and TAKASHI ISHIWATA, Faculty of Information Sciences, Hiroshima City University, Asa-Minami, Hiroshima 731-3194, Japan.

We have generated SiCN in supersonic free jet expansions, and observed the laser induced fluorescence ( LIF ) of the Ã ^{2}Δ – X̃ ^{2}Π transition. We have measured rotationally resolved LIF excitation spectra of the three vibronic bands. Combining with rotational reported transitions^{a}, the rovibronic transitions of the three vibronic bands were analyzed simultaneously, and the spin-orbit constant of the X̃ ^{2}Π state has been determined with precision of the rotational spectroscopy.

aA. J. Apponi, M. C. McCarthy, C. A. Gottlieb, and P. Thaddeus, Astrophys. J. 536, L55 (2000).